20 g of an acid furnishes 0.5 moles if h3o^+ ions in its aqueous solution exactly neutralize 4.9 g of H2So4

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
 H2SO4+2H2O= 2H3O^+ + SO4^2-
no of moles of H2SO4= 4.9/98=1/20.
no. of moles of H3O=1/10;
Amount of H3O ions=19/10=1.9g;
Amount of H3O ions required to react with 20 g of acid=9.5g
Amount of unknown acid = 20x5=100.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)